OpenStudy (anonymous):

first correct answer will get a fan and medal!!! what are the x- and y- intercepts of the line? 2x+3y=-12 a.) x- intercept -6, y-intercept 4 b.) x- intercept -6 y- intercept -4 c.) x-intercept -4 y- intercept 6 d.) x- intercept 4 y-intercept -6

OpenStudy (anonymous):

to find the x intercept, sub in 0 for y 2x + 3y = -12 2x + 3(0) = -12 2x = -12 x = -6 to find the y intercept, sub in 0 for x 2x + 3y = -12 2(0) + 3y = -12 3y = -12 y = - 4 answer is B. Any questions ?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

glad to help :)